# Stable and Accurate Orbital-Free DFT Powered by Machine Learning

### v1

@@ A.6 Density optimization
Figure S.2: Unphysical negative electron densities are admitted by the representation in Eq. 2 but are not a problem in practice. Shown is a histogram of integrated negative densities after density optimization for 200 random QM9 molecules and 6000 optimization steps. While negative density contributions are vanishingly small for STRUCTURES25, a number of densities optimized through the M-OFDFT functional contain significant negative contributions. The number of negative integrated densities below $10^{-8}$ electrons is not depicted but given in parentheses after the respective functional name.
S.3 Robustness with respect to initialization
Zhang et al. [13] require a precise machine-learned guess, “ProjMINAO”, to produce their best results. Their results deteriorate by an order of magnitude when initializing with Hückel [66] densities and become worse still when initializing with a MINAO guess [13, 43, 44]. To gauge the robustness of the STRUCTURES25 functional, we compare density optimizations initialized with our data-driven superposition of atomic densities (dSAD, section A.5) as well as MINAO and Hückel densities. Unlike [13] we refrain from training a model to improve on the initially guessed density.
For the following comparisons, we randomly sample 1000 molecules from our QM9 test set. As a point of reference for the various initial guesses, we report their density error with respect to the ground state in table A.4. Predicted ground state density errors, $\rho^{\mathrm{gs}}-\rho^{\ast}_{\mathrm{guess}}$, from dSAD and MINAO are nearly identical while Hückel fares approximately an order of magnitude worse.
On the same set of molecules, we compare density optimization results starting from the different guesses in Table S.1 and Fig. S.3. Using the identical hyperparameter configuration, both dSAD and MINAO guesses lead to all molecules converging, i.e. achieving gradient norms below $10^{-4}~{}{\mathrm{Ha}}$, within 5000 optimization steps. Hückel initialization is worse, with only 79% of molecules converging with default settings. Increasing the maximum number of iterations to 20k increases the convergence ratio to 85%. Tuning of the momentum to a value of 0.77 pushes the Hückel convergence ratio to 96.7% within 20k iterations. The ability to converge to good solutions for all molecules from both dSAD and MINAO guesses using the same model is a testament to the generality of the STRUCTURES25 functional, as these two initializations differ considerably (cf. Table A.4).
@@
Table S.1: STRUCTURES25 density optimization results with different initializations. Compared are predicted ground state densities from dSAD, MINAO and Hückel guesses on 1000 molecules of the QM9 test set. For the $L_{2}$ density errors per electron, we report the mean. For the number of required iterations to reach convergence we show the median [minimum, maximum] number of iterations. A hyphen “–” indicates lack of convergence within the allowed number of iterations. The first three rows correspond to density optimization configured by the default hyperparameter settings (see section A.6). The last row (Hückel*) shows the results of an additional Hückel run with reduced momentum and greater number of allowed iterations.
@@
Our primary contribution lies in generating training data that is both more diverse and more evenly distributed across the energy landscape as shown in Fig. 3C. To quantify the benefit of this approach, we trained identical models on the QM9 dataset, once using only the standard Kohn-Sham SCF iterations (unperturbed) and once using our perturbed Fock matrix approach. Since the perturbed data has about 1.8 times the number of training labels, we increase the number of epochs for the non-perturbed model trainings to 161. Table S.2 clearly demonstrates the advantages of perturbed data: the resulting model exhibits significantly improved convergence, achieving lower density errors.
@@
We explore the impact of “tensorial” messages [19] in equivariant message passing based on local canonicalization, which allow the communication of non-scalar geometric information between nodes. We evaluate the performance of the standard Graphormer, which uses only scalar messages, against our modified version incorporating tensorial messages, as described in section A.2.
Table A.1 shows results for the QM9 dataset. The additional geometric information improves the model, with the energy error improving slightly and the density error significantly. Networks trained with tensorial messages also showed lower gradient loss during training.

### v2

@@ Acknowledgments
Figure S.5: Unphysical negative electron densities are admitted by the representation in Eq. 2 in the main text, but are not a problem in practice. Shown is a histogram of integrated negative densities after density optimization for 200 random QM9 molecules and 6000 optimization steps. While negative density contributions are vanishingly small for STRUCTURES25, a number of densities optimized through the M-OFDFT functional contain significant negative contributions. The number of negative integrated densities below $10^{-8}$ electrons is not depicted but given in parentheses after the respective functional name.
S.9 Robustness with respect to initialization
Zhang et al. [26] require a precise machine-learned guess, “ProjMINAO”, to produce their best results. Their results deteriorate by an order of magnitude when initializing with Hückel [81] densities and become worse still when initializing with a MINAO guess [26, 55, 56]. To gauge the robustness of the STRUCTURES25 functional, we compare density optimizations initialized with our data-driven superposition of atomic densities (dSAD, section S.5) as well as MINAO and Hückel densities. Unlike prior work [26] we refrain from training a model to improve on the initially guessed density.
For the following comparisons, we randomly sample 1000 molecules from our QM9 test set. As a point of reference for the various initial guesses, we report their density error with respect to the ground state in table S.4. Predicted ground state density errors, $\rho^{\mathrm{gs}}-\rho^{\ast}_{\mathrm{guess}}$, from dSAD and MINAO are nearly identical while Hückel fares approximately an order of magnitude worse.
On the same set of molecules, we compare density optimization results starting from the different guesses in Table S.7 and Fig. S.6. Using the identical hyperparameter configuration, both dSAD and MINAO guesses lead to all molecules converging, i.e. achieving gradient norms below $10^{-4}~{}{\mathrm{Ha}}$, within 5000 optimization steps. Hückel initialization is worse, with only 79% of molecules converging with default settings. Increasing the maximum number of iterations to 20k increases the convergence ratio to 85%. Tuning of the momentum to a value of 0.77 pushes the Hückel convergence ratio to 96.7% within 20k iterations. The ability to converge to good solutions for all molecules from both dSAD and MINAO guesses using the same model is a testament to the generality of the STRUCTURES25 functional, as these two initializations differ considerably (cf. Table S.4).
@@
Table S.7: STRUCTURES25 density optimization results with different initializations. Compared are predicted ground state densities from dSAD, MINAO and Hückel guesses on 1000 molecules of the QM9 test set. For the $L_{2}$ density errors per electron, we report the mean. For the number of required iterations to reach convergence we show the median [minimum, maximum] number of iterations. A hyphen “–” indicates lack of convergence within the allowed number of iterations. The first three rows correspond to density optimization configured by the default hyperparameter settings (see section S.6). The last row (Hückel*) shows the results of an additional Hückel run with reduced momentum and greater number of allowed iterations.
@@
Our primary contribution lies in generating training data that is both more diverse and more evenly distributed across the energy landscape as shown in Fig. 3C. To quantify the benefit of this approach, we trained identical models on the QM9 dataset, once using only the standard Kohn-Sham SCF iterations (unperturbed) and once using our perturbed Fock matrix approach. Since the perturbed data has about 1.8 times the number of training labels, we increase the number of epochs for the non-perturbed model trainings to 161. Table S.8 and Fig. S.7 clearly demonstrate the advantages of perturbed data: the resulting model exhibits significantly improved convergence and achieves lower density errors.
@@
Figure S.7: Perturbation of the external potential for training data improves convergence in density optimization. In density optimization, the model trained on conventional data generated without perturbing the external potential (purple) often does not converge while the model trained with data generated with our perturbation scheme consistently converges within a few hundred iterations (orange).
@@
We explore the impact of “tensorial” messages [31] in equivariant message passing based on local canonicalization, which allow the communication of non-scalar geometric information between nodes. We evaluate the performance of the standard Graphormer, which uses only scalar messages, against our modified version incorporating tensorial messages, as described in section S.2.
Table S.1 shows results for the QM9 dataset. The additional geometric information improves the model, with the energy error improving slightly and the density error significantly. Networks trained with tensorial messages also showed lower gradient loss during training.
@@
S.11 Runtime comparison of KS-DFT and OF-DFT computations
Figure S.8 compares the runtimes of OF-DFT as embodied by STRUCTURES25 and KS-DFT as implemented in GPU4PySCF [89, 90], which can exploit the parallelism of the identical hardware. All computations were run on one Nvidia A100 GPU and 8 cores of an AMD EPYC 7452 processor. For the largest molecule of our QMUGS test set, STRUCTURES25 reaches a speedup of almost an order of magnitude. Importantly for future calculations on larger (e.g., biomolecular) systems, it shows favorable scaling with the size of the system.
Figure S.8: Comparison of runtimes of KS-DFT and OF-DFT computations.